OpenStudy (anonymous):

The diameters of two cylinders are in the ratio of 2 : 1 and their volumes are equal. The ratio of their heights will be. (a) 1 : 6 (b) 1 : 2 (c) 1 : 4 (d) 3 : 4

OpenStudy (anonymous):

plz give the answer...

OpenStudy (anonymous):

The volume of a cylinder is \[V = \pi r^2*h = {\pi \over 4} D^2 *h\]Since the volumes are equal\[{\pi \over 4} (2D)^2*h_1 = {\pi \over 4}D^2*h_2\]Solving the ration h1/h2 gives

OpenStudy (anonymous):

Oops! Let me finish that

OpenStudy (anonymous):

\[{h_1 \over h_2} = {(2D^2) \over D^2}\]

OpenStudy (anonymous):

options plz

OpenStudy (anonymous):

You're going to have to figure that out on your own. I don't mind coaching you through a problem, but I don't like giving out answers. You won't learn it if I give you answers.

OpenStudy (anonymous):

plz..plz...plz....sir